A). Determine the following limits (if they exist)

1. Lim X approaches -5 x^2+x-20/3(X+5)

2. Lim t approaches 0 sin5t/t^2+4t

3. Lim approaches -infinity 3-|X|/2|X|+1

4. Lim X approaches 0 2x/3-√x+9

5. Lim X approaches +infinity 2x+x^2+1/1-x+2x^2


1
Expert's answer
2021-08-17T10:14:11-0400

"\displaystyle
1.\\
\begin{aligned}
\lim_{x \to -5} \frac{x^2+x-20}{3(x+5)} &= \lim_{x \to -5} \frac{(x + 5)(x - 4)}{3(x+5)}
\\&= \lim_{x \to -5} \frac{x - 4}{3}
\\&= \frac{-9}{3} = -3
\end{aligned}\\

2.\\
\begin{aligned}
 \lim_{x \to 0} \frac{\sin(5t)}{t^2+4t} &= \lim_{x \to 0} \frac{5\cos(5t)}{2t + 4}
\\&= \frac{5}{4}
\end{aligned}\\

3.\\
\begin{aligned}
\lim_{x \to -\infty} \frac{3-|x|}{2|x|+1} &= \lim_{x \to -\infty} \frac{\frac{3}{|x|} - 1}{2+\frac{1}{|x|}} 
\\&= \frac{-1}{2}
\end{aligned}\\


4.\\
\begin{aligned}
\lim_{x \to 0} \frac{2x}{3 - \sqrt{x} + 9} &= \frac{2(0)}{3 - \sqrt{0} + 9} 
\\&= \frac{0}{12} = 0
\end{aligned}\\


5.\\
\begin{aligned}
\lim_{x \to \infty} \frac{2x+x^2+1}{1-x+2x^2}
&= \lim_{x \to \infty} \frac{2/x + 1 + 1/x^2}{1/x^2 - 1/x +2}
\\&= \frac{1 + 0 + 0}{2 - 0 + 0} = \frac{1}{2}
\end{aligned}"